OpenStudy (anonymous):

factor the polynomial expression 9x^2-16 medal and fan

OpenStudy (ivettef365):

that is (a^2-b^2) = (a+b)(a-b)

OpenStudy (akashdeepdeb):

Write it in the form of a^2 - b^2 = (a+b)(a-b) Like (3x)^2 - 4^2 Now use the identity! Getting me? :)

OpenStudy (anonymous):

thanks ;)

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

(3x+4)(x-4)

OpenStudy (akashdeepdeb):

Yeah just put a 3 over there and you're done! :D

OpenStudy (anonymous):

(3x+4)(3x-4)? like that

OpenStudy (ivettef365):

perfect

OpenStudy (akashdeepdeb):

Exactly! :D

OpenStudy (anonymous):

thanks

OpenStudy (akashdeepdeb):

:)